# Efficient Depth Estimation for Unstable Stereo Camera Systems on AR Glasses

### v1

@@ Abstract
Stereo depth estimation is a fundamental component in augmented reality (AR) applications. Although AR applications require very low latency for their real-time applications, traditional depth estimation models often rely on time-consuming preprocessing steps such as rectification to achieve high accuracy. Also, non standard ML operator based algorithms such as cost volume also require significant latency, which is aggravated on compute resource-constrained mobile platforms. Therefore, we develop hardware-friendly alternatives to the costly cost volume and preprocessing and design two new models based on them, MultiHeadDepth and HomoDepth. Our approaches for cost volume is replacing it with a new group-pointwise convolution-based operator and approximation of consine similarity based on layernorm and dot product. For online stereo rectification (preprocessing), we introduce homograhy matrix prediction network with a rectification positional encoding (RPE), which delivers both low latency and robustness to unrectified images, which eliminates the needs for preprocessing. Our MultiHeadDepth, which includes optimized cost volume, provides 11.8-30.3% improvements in accuracy and 22.9-25.2% reduction in latency compared to a state-of-the-art depth estimation model for AR glasses from industry. Our HomoDepth, which includes optimized preprocessing (Homograhpy + RPE) upon MultiHeadDepth, can process unrectified images and reduce the end-to-end latency by 44.5%. We adopt a multi-task learning framework to handle misaligned stereo inputs on HomoDepth, which reduces theAbsRel error by 10.0-24.3%. The results demonstrate the efficacy of our approaches in achieving both high model performance with low latency, which makes a step forward toward practical depth estimation on future AR devices.
@@ 1 Introduction
Depth estimation serves as a foundational component in augmented and virtual reality (AR/VR) [16] with many downstream algorithms, which include novel-view rendering [14, 28], occlusion reasoning [17], world locking for AR object placement [18], and determining the scale of AR objects [3]. In the AR domain, stereo depth estimation is often deployed [28] rather than mono depth estimation due to its superior accuracy and the ease of deploying stereo cameras on the each side of the AR glasses frame. One key objective for depth estimation models targeting AR glasses is the low latency, in addition to the good performance, because many AR use cases are realtime applications [16]. However, since AR glasses are in a compute resource-constrained wearable form factor, enabling desired latency (less than 100 ms on device) is not trivial.
@@
One challenge towards the latency optimization originates from the preprocessing, which is a indispensable step for achieving high model performance in practical applications. Examples include camera distortion and alignment using camera intrinsic and extrinsic. In addition to image preprocessing, depth estimation models often involve traditional algorithms such as cost volume [7, 28]. We present their significance in latency in Fig. 1, which shows that preprocessing accounts for 30.2% and cost volume accounts for 29.3% of the total latency of a state-of-the-art model, Argos [28]. Note that preprocessing latency can be more dominant when camera intrinsic and extrinsic parameters are unknown, which results in 200 to 2000 ms latency to solve for the extrinsic parameters and subsequently process the images [15, 30]. This aggravates the long preprocessing latency challenge further.
@@
Therefore, to reduce depth estimation model latency for AR glasses, we propose new methodologies that significantly reduce the preprocessing (online stereo rectification not required) and cost volume latency, as shown in Fig. 1. Our approach for the cost volume is (1) to replace traditional algorithm with group-pointwise convolutions, which are highly optimized in hardware and compiler and (2) adopt an efficient approximation of consine similarity using layernorm and dot product. For preprocessing (stereo matching), we adopt homography matrix-based approximation and estimate homography using a head attached to the depth estimation model, which allows to utilize homography matrix-based approximation with dynamically varying extrinsic parameters in unstable AR glasses platform or no access to camera extrinsic parameters.
@@
Since our methodologies are complementary to existing depth estimation models, we augment our methodologies on the Argos [28] and develop two versions of new models, MultiHeadDepth and HomoDepth, to show the effectiveness of our methodologies. As presented in Fig. 1, MultiHeadDepth focuses on cost volume optimization, which reduces AbsRel by 7% and inference latency by 25% compared to the original cost volume. HomoDepth targets scenarios that require stereo matching preprocessing, which can directly accept unrectified images as input, eliminating the needs for preprocessing step. HomoDepth not only provides high-quality outputs on unrectified images as presented in Tab. 4, but also significantly reduces the end-to-end latency by 43%. Our evaluation includes ADT [20] dataset collected by real research AR glasses by Meta, Aria [5], which demonstrates the effectiveness of our approach on realistic AR glasses platform.
@@
  • We develop multi-head cost volume block that replaces costly cost volume blocks in previous depth estimation models. Our new block provides significantly lower latency as well as higher accuracy compared to the original cost volume.
  • We introduce the homograpghy of stereo images to reveal their position relationship to enable to accept unrectified images without preprocessing. We merge a small homograpghy estimation head within the depth estimation network, which significantly reduces the latency compared to the preprocessing-based approach.
  • We augment our homography estimation head with 2D rectification position encoding, which helps translate relative positional information from homograpghy matrix to emdeding coding format. It enables the neural network to effectively understand relative position information, which plays a key role in eliminating the need for rectification preprocessing.
Figure 1: The latency breakdown analysis of a state-of-the-art model Argos [28] and ours, MultiHeadDepth & HomoDepth on Intel Core(TM) i7-12700H CPU (laptop). The ”RPE” refers to the 2D rectification position encoding process. ”Others” refers to all the other part of the neural network excluding cost volume blocks, such as Conv, Norm, FC, and ReLU6.
@@ 2 Related Work
Argos [28] extended Tiefenrausch monocular network [14] to implement stereo depth estimation. Argos is optimized to minimize the model size and computational complexity. Argos Network consists of encoder, cost volume, and a decoder, which requires preprocesing if the input images are not rectified or distorted. Argos was designed for practical AR applications by Meta, and it delivers state-of-the-art performance in the stereo depth estimation on AR glasses domain.
DynamicStereo [11] is another lightweight design in the parameter size with low latency as well, but Argos delivers superior accuracy and latency.
@@
Stereo Image Preprocessing. The preprocessing of stereo images consists of two parts: Calibration and Rectification. Calibration is required for a raw image captured by the camera, which removes distortion and artifacts in the raw image caused by the imaging system. Sometimes it crops images to remove the black margins around the images. Calibration depends on the intrinsic parameters of camera systems. Generally, the intrinsic of a camera is stable and provided by the manufacturer. Calibration can be completed with just a few matrix computations, requiring only a few dozen milliseconds of hardware.
@@
Rectification is a major challenge in stereo vision systems. For an object in the real world, its projection onto the two image planes should lie on the epipolar lines, satisfying the epipolar constraint [6]. If there is a misalignment between two cameras, it is necessary to rectify the pair of images so that corresponding points lie on the same horizontal line. The stereo cameras are mounted on both sides of AR glasses. The glasses may undergo significant bending ($>10^{\circ}$) because of the soft material. Additionally, this issue is further exacerbated by variations in users’ head sizes and temperatures. [28].
Rectification requires the relative position information of the two cameras. In previous research, taking MobiDepth [30] as an example, it assumed that the relative positions of the cameras had been obtained before the inference phase. It can only process images based on the known position relationship. This is considered as offline rectification. However, in practical applications, the camera positions on AR glasses need to be determined based on a pair of arbitrarily captured images, which results in online rectification.
@@
Karaev et al. [11] set image saturation to a value sampled uniformly between 0 and 1.4 during the model training phase to make the model more robust for stereo images misalignment. But it lacks good interpolation for robustness. Wang et al. proposed a fast online rectification [28]. However, there remains a 15% to 23% chance of failing to rectify the stereo images.
Every stereo-matching algorithm includes a mechanism for evaluating the similarity of image locations. Matching cost is calculated for every pixel over the range of examined disparities, and it’s a commonly used measurement [9].
@@
Kendall et al. [12] adopted deep unary features to compute the stereo matching cost by forming a cost volume. They form a cost volume of dimensionality $height\times weight\times(max\_disparity+1)\times features$ and pack it into a 4D volume. To construct the 4D volume, the algorithm iteratively move one of the images/features horizontally and compare the matching cost in each disparity, as described in Fig. 2(a). The cost volume information is crucial for a neural network to understand the differences between the two feature maps. Although proven effective, the cost volume has a disadvantage in the computational cost that originates from cosine similarity computation described below:
$D_{cos}(\boldsymbol{a},\boldsymbol{b})=\frac{\boldsymbol{a}\cdot\boldsymbol{b}}{||\boldsymbol{a}||||\boldsymbol{b}||}$ (1)
@@
where $\boldsymbol{a}$ and $\boldsymbol{b}$ are the vectors of pixels. The numerator is matrix multiplication-optimized hardware- and compiler-friendly since the dot product computation in a batch is matrix multiplication. However, the denominator requires to compute norms pixel by pixel, which is not very friendly to hardware highly optimized for matrix multiplication.
@@ 3 MultiHeadDepth Model
Based our observation on the cost volume, we optimize it by approximating cosine similarity with a set of hardware-friendly operations and introduce a new hardware-friendly block, multi-head cost volume, to replace the overall cost volume algorithm.
@@ 3.1 Approximating Cosine Similarity
The cosine similarity calculates the angle between two vectors, which indicates the similarity of the vectors. Since the denominator needs to be optimized as we discussed in Sec. 2, we focus on the denominator. Our approach is replacing the division by denominator in  Eq. 1 with 2D layer normalization (layer norm) since this approach reduces the computational complexity and achieves normalization effect (i.e., the elements within a pixel vector have a mean of zero and a standard deviation of one). Maintaining the numerator (dot product) as is in the modified distance metric, the computational complexity is reduced, as discussed in Sec. 7.1. Although the new metric is not mathematically equivalent as the original cosine similarity, but our evaluation results indicate that the metric is effective.
Another benefit of the layer norm is that it provides a buffer for the cost volume block, enabling it to accept additional encoded information. The pixel of input activation is normalized so that it can be merged with other encodings. For instance, it can handle positional encoding, which will be introduced in Sec. 4.4.
@@ 3.2 Multi-head Cost Volume
Figure 2: The comparison between the original and multi-head cost volume
Inspired by multi-head attention mechanism [26], we introduce the multi-head attention and dot scale into the cost volume block since separating the input activation in multiple heads can enable more perceptions of the network. The original cost volume calculates the cosine similarity of pixel vectors of two input activations, and the output is scalar, as shown in Fig. 2(a). The multi-head cost volume computes dot products to produce numerous scalars, forming new pixel vectors. Subsequently, we apply point-wise convolution to generate an output that matches the size of the original cost volume, as Fig. 2(b) shows. Overall, these operations can be executed just by group-wise Conv, which is highly optimized in modern GPUs. The pseudo-code of multi-head cost volume is shown as Algorithm 1.
@@
$left:[NCHW]$
$right:[NCHW]$
$max\_disparity:int$
$head\_num:int$triangleright number of heads, can divide C
@@
    $right\_shifted=\text{roll}(right\_norm,i)$
@@
The benefits of multi-head cost volume can be summarized with (1) less computational complexity, (2) replacing the computation with highly-optimized group-wise Conv, and (3) potential accuracy improvement by providing more perceptions of the input activations
@@ 3.3 Structure of MultiHeadDepth
Applying our approaches discussed in Sec. 3.1 and Sec. 3.2 to Argos [28], we design MultiHeadDepth, which is illustrated in Fig. 3. All settings for the inverted residual blocks [23] in Argos are retained to highlight the efficacy of our Multi-head Cost Volume in an ablation study. Note that our approach can be applied to other models with cost volume as well.
Figure 3: The structure of MultiHeadDepth. The dashed lines indicate that the input activations from the left image are passed to the decoders. The input example is from ADT dataset
@@ 4 HomoDepth Model
Figure 4: Structure of HomoDepth. The orange connections are only valid in training phase; the green connection is only valid in inference phase. The input example is from DTU dataset.
We discuss how we can augment MultiHeadDepth to include opitmizations for preprocessing process, which consists of the introduction of 3D projection and homographic matrix with homography estimation head.
@@
The main functionality of cost volume is stereo-matching, which evaluates the similarity of left and right inputs with different offsets to determine the possible disparity. Because the algorithm only allows horizontal offsets, it can only match patterns on a common horizontal line. Therefore, preprocessing is required for the input images for stereo cameras without good alignments.
We will discuss depth estimation from a broader perspective. We focus on the projection relationship of a world point in the image planes of a stereo camera system. Here, we pose the fundamental question: Given that the image points $q_{l}$ and $q_{r}$ are derived from the same world point $Q$, what is the relationship between $q_{l}$ and $q_{r}$?
Suppose the distance from $Q$ to the left and right image planes are $d_{l},d_{r}$ respectively, intrinsic parameters of left and right cameras are $K_{l},K_{r}$ respectively, and the extrinsic parameters of left and right cameras are $M_{l},M_{r}$ respectively. We derive the relationship between $q_{l}$ and $Q$, $q_{r}$ and $Q$,
@@
where $H_{l\rightarrow r}$ is the plane homography matrix [8] and its size is $3\times 3$. $H_{l\rightarrow r}$ converts $q_{l}$ to $q_{r}$. That is to say, knowing the homographic matrix and distance information, it is easy to represent the position relation between the content in stereo images.
Unfortunately, distance information is unknown in the depth estimation task since it is the eventual goal of the task. However, we can leverage the following properties: ${d_{l}}/{d_{r}}\approx 1$ when $Q$ is on the central plane of the cameras or far from the camera system. Note that most imaged objects maintain a certain distance from the imaging system, in practice.
Thus, homography matrix can approximately represent the positional relationship of stereo image planes, which is helpful for rectification. Homography is widely used in stereo vision applications, and for example, MVSNet [29] has demonstrated the effectiveness of homography.

### v2

@@ Abstract
Stereo depth estimation is a fundamental component in augmented reality (AR), which requires low latency for real-time processing. However, preprocessing such as rectification and non-ML computations such as cost volume require significant amount of latency exceeding that of an ML model itself, which hinders the real-time processing required by AR. Therefore, we develop alternative approaches to the rectification and cost volume that consider ML acceleration (GPU and NPUs) in recent hardware. For pre-processing, we eliminate it by introducing homography matrix prediction network with a rectification positional encoding (RPE), which delivers both low latency and robustness to unrectified images. For cost volume, we replace it with a group-pointwise convolution-based operator and approximation of cosine similarity based on layernorm and dot product.
Based on our approaches, we develop MultiHeadDepth (replacing cost volume) and HomoDepth (MultiHeadDepth + removing pre-processing) models. MultiHeadDepth provides 11.8-30.3% improvements in accuracy and 22.9-25.2% reduction in latency compared to a state-of-the-art depth estimation model for AR glasses from industry. HomoDepth, which can directly process unrectified images, reduces the end-to-end latency by 44.5%. We also introduce a multi-task learning method to handle misaligned stereo inputs on HomoDepth, which reduces the AbsRel error by 10.0-24.3%. The overall results demonstrate the efficacy of our approaches, which not only reduce the inference latency but also improve the model performance. Our code is available at https://github.com/UCI-ISA-Lab/MultiHeadDepth-HomoDepth
@@ 1 Introduction
Depth estimation serves as a foundational component in augmented and virtual reality (AR/VR) [18] with many downstream algorithms, which include novel-view rendering [16, 31], occlusion reasoning [19], world locking for AR object placement [20], and determining the scale of AR objects [3]. In the AR domain, stereo depth estimation is often deployed [31] rather than mono depth estimation due to its superior accuracy and the ease of deploying stereo cameras on the each side of AR glasses frames. Due to the realtime nature of AR applications [18], one key objective for depth estimation models for AR is low latency, in addition to good performance. However, since AR glasses are in a compute resource-constrained wearable form factor, enabling desired latency (less than 100 ms on device) is not trivial.
@@
One challenge towards the latency optimization originates from the preprocessing, which is a indispensable step for achieving high model performance in practical applications. Examples include camera calibration and rectification using camera intrinsic and extrinsic. In addition to preprocessing, traditional algorithms embedded into depth estimation models (e.g, cost volume [8, 31]) also imposes another challenge to the latency optimization. We present their significance in latency in Fig. 1, which shows that preprocessing accounts for 30.2% and cost volume accounts for 29.3% of the total latency of a state-of-the-art (SOTA) model, Argos [31]. Note that preprocessing latency can be more dominant when camera intrinsic and extrinsic parameters are unknown, which results in 200 to 2000 ms latency to solve for the extrinsic parameters and subsequently process the images [17, 34]. This aggravates the long preprocessing latency challenge further.
@@
Therefore, to reduce depth estimation model latency for AR glasses, we propose new methodologies that significantly reduce the preprocessing (online stereo rectification not required) and cost volume latency, as shown in Fig. 1. Our approach for the cost volume is (1) to replace traditional algorithm with group-pointwise convolutions, which are highly optimized in hardware and compilers and (2) adopt an efficient approximation of cosine similarity using layernorm and dot product. For preprocessing (stereo matching), we adopt homography matrix-based approximation and estimate homography using a head attached to the depth estimation model, which allows to utilize homography matrix-based approximation with dynamically varying extrinsic parameters in unstable AR glasses platform or no access to camera extrinsic parameters.
@@
Since our methodologies are complementary to existing depth estimation models, we augment our methodologies on the Argos [31] and develop two new models, MultiHeadDepth and HomoDepth. As presented in Fig. 1, MultiHeadDepth focuses on cost volume optimization, which reduces inference latency by 25% compared to the original cost volume. HomoDepth targets scenarios that require stereo matching preprocessing, which can directly accept unrectified images as input, eliminating the needs for preprocessing. HomoDepth not only provides high-quality outputs on unrectified images as presented in Tab. 5, but also significantly reduces the end-to-end latency by 43%. Our evaluation includes ADT [22] dataset collected by real research AR glasses by Meta, Aria [6], which demonstrates the effectiveness of our approach on realistic AR glasses platform.
@@
  • We develop MultiheadCostVolume block that replaces cost volume in a previous depth estimation model. Our new block provides significantly lower latency as well as higher accuracy.
  • We introduce the homography of stereo images to reveal their position relationship to enable to accept unrectified images without preprocessing. We merge a small homography estimation head within the depth estimation network, which significantly reduces the latency compared to the preprocessing-based approach.
  • We augment our homography estimation head with 2D rectification position encoding, which helps translate relative positional information from homograpghy matrix to the 2D rectification position encoding format. It enables the neural network to effectively understand relative position information, which plays a key role in eliminating the need for rectification preprocessing.
Figure 1: The latency breakdown analysis of a SOTA model Argos [31] and ours, MultiHeadDepth & HomoDepth on Intel i7-12700H laptop CPU. The ”RPE” refers to the 2D rectification position encoding process. ”Others” refers to all the other parts of the neural network excluding cost volume blocks, such as Conv, Norm, FC, and ReLU6.
@@ 2 Related Work
Argos [31] extended Tiefenrausch monocular network [16] to implement stereo depth estimation. Argos is optimized to minimize the model size and computational complexity. Argos consists of an encoder, cost volume, and a decoder, which requires preprocesing if the input images are distorted or not rectified. Argos was designed for practical AR applications by Meta, and it delivers SOTA performance in the stereo depth estimation on AR glasses domain.
DynamicStereo [13] is a stereo depth estimation model designed for AR and mobile devices with a focus on efficiency, but Argos delivers superior accuracy and latency for just one-shot estimation.
@@
Stereo Image Preprocessing. The stereo image preprocessing consists of calibration and rectification. Calibration is required for a raw image captured by the camera, which removes distortion and artifacts in the raw image. Calibration depends on the intrinsic parameters of camera systems. Generally, the intrinsic of a camera is stable and provided by manufacturers. Calibration can be completed with just a few matrix computations, within a few dozen milliseconds.
@@
Rectification is a major challenge in stereo vision systems. For an object in the real world, its projection onto two image planes should lie on the epipolar lines, satisfying the epipolar constraint [7]. If the two cameras are misaligned, it is necessary to rectify the pair of images so that corresponding points lie on the same horizontal lines. The stereo cameras are mounted on both sides of AR glasses. The glasses may undergo significant bending ($>10^{\circ}$) because of the soft material [31]. This issue can be exacerbated by variations in users’ head sizes and temperatures.
@@
Rectification requires the relative position information of two cameras. MobiDepth [34] assumed that the relative positions of the cameras had been obtained before the inference. However, practically, the camera positions on AR glasses need to be determined based on a pair of arbitrarily captured images, which requires online rectification unlike MobiDepth’s offline method. Karaev et al. [13] set image saturation to a value sampled uniformly between 0 and 1.4 during the model training phase to make the model more robust for stereo images misalignment. But it lacks good interpolation for robustness. Wang et al. proposed a fast online rectification [31]. However, there remains a 15% to 23% chance of failing to rectify the stereo images.
@@
Every stereo-matching algorithm includes a mechanism for evaluating the similarity of image locations. Matching cost is calculated for every pixel over the range of examined disparities, and it’s a commonly used measurement [11]. Kendall et al. [14] adopted deep unary features to compute the stereo matching cost by forming a cost volume. They form a cost volume of dimensionality $height\times weight\times max\_disparity\times features$ and pack it into a 4D volume. To construct the 4D volume, the algorithm iteratively move one of the images/features horizontally and compare the matching cost in each disparity, as described in Fig. 3(a). The cost volume information is crucial for a neural network to understand the differences between the two feature maps. Although proven effective, the cost volume has a disadvantage in the computational efficiency from its cosine similarity computation described below:
$D_{cos}(\boldsymbol{a},\boldsymbol{b})=\boldsymbol{a}\cdot\boldsymbol{b}~{}/~{}|\boldsymbol{a}||\boldsymbol{b}|$ (1)
@@
where $\boldsymbol{a}$ and $\boldsymbol{b}$ are the vector of pixels. The numerator is matrix multiplication-optimized hardware- and compiler-friendly since the dot product computation in a batch is a matrix multiplication. However, the denominator requires to compute norms pixel by pixel, which is not friendly to hardware highly optimized for matrix multiplication. Therefore, we first focus on the optimization of cost volume, which we discuss next.
@@ 3 MultiHeadDepth Model
Our approach to optimize cost volume is to replace it by a hardware-friendly approximation. We discuss our approach in detail and our solution, multi-head cost volume, that implements our approach.
Figure 2: Similarity estimation methodology Comparison. Each represents a similarity map between left and right features after applying roll operation (offset: 10) to the input stereo images. The maps are average results across the entire dataset of Sceneflow [21], after rescaling to [0,1] range. The strips on the left side of maps are caused by roll and they are part of the maps.
@@ 3.1 Approximating Cosine Similarity
As discussed in Sec. 2, the denominator of cosine similarity (computing two norms and multiplication of them) and the corresponding division are the prime optimization targets. The goal of those operations is to normalize individual vectors to bound the results in the [-1,1] range. Based on the observation, we approximate the normalization for individual vectors using 2D layer normalization. This approach significantly reduces the number of costly normalization computations, as discussed in Appendix A. We still keep the numerator of cosine similarity as a dot product operator after the normalization. We term our approach as LND (LayerNorm combined with Dot product). LND successfully approximates the cosine similarity, as shown by the results in Fig. 2. Another benefit of the layer norm is that it provides a buffer for the cost volume, which can accept additional encoded information. That is, each normalized pixel can be merged with other encodings. For instance, it can handle positional encoding, which we discuss in Sec. 4.4.
To further enhance the model performance, we also augment the LND with a weight layer after the LND, which help adjust the similarity approximation. Our evaluation results with better model performance and latency show the effectiveness of our approach that combines an efficient approximation and adaptive weight layers.
@@ 3.2 Multi-head Cost Volume
Figure 3: The comparison between the original cost volume and our approach, MultiheadCostVolume
As another enhancement, we introduce the multi-head attention and dot scale [29] into our cost volume since separating the input activation in multiple heads enable more perceptions of the network. The approach also leads to a hardware-friendly operator, group-wise dot product followed by a point-wise convolution, as shown in Fig. 3 (b). This replaces the costly cosine similarity depicted in Fig. 3 (a), while providing better model performance. We describe the multi-head cost volume algorithm in a pseudo-code in Algorithm 1.
Figure 4: The structure of MultiHeadDepth. The dashed lines indicate that the input activations from the left image are passed to the decoders. The input example is from ADT dataset.
@@
$left:[NCHW],~{}~{}right:[NCHW],~{}~{}max\_disparity:int,$
$head\_num:int$triangleright number of heads, can divide C evenly
@@
    $right\_shifted=\text{roll}(right\_norm,i)$triangleright roll: right-shifts the image by i pixels, and the right-most i columns move to the left
@@
The benefits of multi-head cost volume can be summarized with (1) less computational complexity, (2) replacing the computation with highly-optimized group-pointwise Conv, and (3) potential accuracy improvement by providing more perceptions of the input activations.
@@ 3.3 Structure of MultiHeadDepth
Applying our approaches discussed in Sec. 3.1 and Sec. 3.2 to Argos [31], we design MultiHeadDepth, which is illustrated in Fig. 4. All settings for the inverted residual blocks [25] in Argos are retained to highlight the efficacy of our Multi-head Cost Volume in an ablation study. Note that we apply our approach to Argos since it is the state-of-the-art today; our approach can be applied to any other depth estimation models with cost volume as well.
Upon this model, MultiHeadDepth, we add an optimization targeting preprocessing, which we discuss next.
@@ 4 HomoDepth Model
Figure 5: Structure of HomoDepth. The input example is from DTU dataset.
We discuss how we augment MultiHeadDepth with an optimization for preprocessing. We first introduce the 3D projection and homographic matrix and discuss how we estimate it using a homography estimation head.
@@
The main functionality of cost volume is stereo-matching, which evaluates the similarity of left and right inputs with different offsets to determine the possible disparity. Because the algorithm only allows horizontal offsets, it can only match patterns on a common horizontal line. Therefore, preprocessing is required for the input images for stereo cameras without good alignments. We discuss depth estimation from a broader perspective, focusing on the projection relationship of a world point in the image planes of a stereo camera system. Aligned with that, we pose a fundamental question: Given that the image points $q_{l}$ and $q_{r}$ are derived from the same world point $Q$, what is the relationship between $q_{l}$ and $q_{r}$?
Suppose the distance from $Q$ to the left and right image planes, intrinsic parameters, and extrinsic parameters are ($d_{l},d_{r}$), ($K_{l},K_{r}$), and ($M_{l},M_{r}$), respectively. Then the the relationship between $q_{l}$ and $Q$, $q_{r}$ and $Q$ is as follows:
@@
where $H_{l\rightarrow r}$ is the $3\times 3$ plane homography matrix [10], which converts $q_{l}$ to $q_{r}$. That is, it is easy to represent the positional relation between the content in stereo images if we know the homographic matrix and distance information.
Unfortunately, distance information is unknown in the depth estimation task since it is the eventual goal of the task. However, we can still leverage the following property: ${d_{l}}/{d_{r}}\approx 1$ when $Q$ is on the central plane of the cameras or far from the cameras. Note that most imaged objects maintain a certain distance from the imaging system, in practice. Thus, homography matrix can approximately represent the positional relationship of stereo image planes, which is helpful for rectification. Homography is widely used in stereo vision applications, and for example, MVSNet [33] has demonstrated the effectiveness of homography.